# Evidence for Dynamical Dark Matter

### v1

@@ A. Perturbation formula
Figure 1: One-dimensional and two-dimensional posterior distributions of four free parameters in the extended DDM model. The blue dashed lines are $\omega_{dm0}=\omega_{dma}=c^{2}_{s,eff}=\sigma_{dm}=0$, while the red dotted line denotes the linear relation $\omega_{dma}=-\omega_{dm0}$.
@@ C. The effect of curvature on the dynamical dark sector
The cosmic curvature $\Omega_{K}$ plays an important role in depicting the background evolution and structure formation of the universe. Here we study its impact on the coexistence of DDM and DDE in the DDME scenario. Overall, we find that $\Omega_{K}$ hardly affect our conclusions. Interestingly, CMB gives $\omega_{dma}=0.164^{+0.093}_{-0.057}$, which is a $2.53\,\sigma$ evidence for the redshift evolution of DM. The addition of $\Omega_{K}$ increases the DDM evidence from $1.82\,\sigma$ to $2.53\,\sigma$. CD, CP and CDP provides a $1.93\,\sigma$, $2.40\,\sigma$ and $1.31\,\sigma$ evidence of DDM, respectively. Furthermore, we find that C, CD, CP and CDP exhibit a $\sim 2\,\sigma$, $2.37\,\sigma$, beyond $2\,\sigma$, and $2.40\,\sigma$ signal of DDE, respectively. It is worth noting that in the planes of $\omega_{dm0}-\omega_{0}$, $\omega_{dm0}-\omega_{a}$, $\omega_{dma}-\omega_{0}$ and $\omega_{dma}-\omega_{a}$, C, CD, CP and CDP still support a beyond $2\,\sigma$ evidence of the coexistence of DDM and DDE in the dark sector of the universe. Additionally, the $\sim 2\,\sigma$ preference of a closed universe by Planck CMB data could be alleviated when simultaneously considering the DDM and DDE.
@@
D. Destiny+
In Wang:2025owe , based on the latest cosmological observations, we have proposed the new fate of the universe called the “Big Stall” in the CPL DDE model. In this work, when considering the coexistence of DDM and DDE, we propose a new density of the universe, i.e., the “Super Rip” in the DDME model. They both support that the universe is ultimately dominated by DM rather than DE. In Figs.3 and 4, we show the $1\,\sigma-5\,\sigma$ regions of background quantities $\Omega_{m}(z)$, $q(z)$, $E(z)$ and $\omega_{dm}(z)$ from different datasets in the DDM, SDDM and DDME scenarios. The details of their properties and the Super Rip have been specified in the main text.
@@ D. Destiny+
E. Special example: constraints on $\omega_{dm0}$
@@
F. Reconstructing $\omega_{dm}(z)$ with Guassian process
@@
where the prime denotes the derivative with respect to the redshift $z$. The background quantities $D^{\prime}$ and $D^{\prime\prime}$ in Eq. (4) can be easily obtained using the GP reconstructions. In our analysis, for simplicity, we consider a flat universe, namely $\Omega_{k}=0$ and assume $H_{0}=73.04\pm 1.04$ km s-1 Mpc-1 Riess:2021jrx . We also take $\Omega_{m}=0.3153\pm 0.0073$ and $\Omega_{b}=0.04932\pm 0.00052$ Planck:2018vyg1 . For DESI BAO data, assuming $r_{d}=147.09\pm 0.26$ Mpc, we transform $D_{M}(z)/r_{d}$ and $D_{H}(z)/r_{d}$ to $D$ and $D^{\prime}$, respectively. For CC, we directly transform $H(z)$ to $D^{\prime}$, while transforming $D_{L}(z)$ from SN and $\mathcal{R}(z_{c})=1.7488\pm 0.0074$ Planck:2015bue ($z_{c}=1089.0$ is the redshift of recombination) from CMB to $D$. Same as Seikel:2012uu , we use the so-called Matérn ($\nu=9/2$) covariance function in GP reconstructions. Based on Eq.4 and different kindd of datasets, we use a modified version of the package GaPP to carry our the GP reconstructions.
@@
G. Comparison of DESI DR1 and DR2
In Figs.7 and 8, we demonstrate the robustness of our conclusions and show the improvements of the constraining power of DESI DR2 BAO measurements relative to DESI DR1 data in the SDDM and DDME models. One can easily find that the constraints from DESI DR1 and DR2 data on evolution of DDM is robust and that the high-precision DESI DR2 BAO data significantly compresses the parameter space of DDE.

### v2

@@ A. Perturbation formula
Figure 1: One-dimensional and two-dimensional posterior distributions of four free parameters from the CMB dataset in the extended DDM model. The blue dashed lines are $\omega_{dm0}=\omega_{dma}=c^{2}_{s,eff}=\sigma_{dm}=0$, while the red dotted line denotes the linear relation $\omega_{dma}=-\omega_{dm0}$.
@@
Figure 3: One-dimensional posterior distributions of the parameter $\Sigma m_{\nu}$ from different datasets in the $\Lambda\omega$DM, SDDM, DDM and DDME models.
@@ B. The four-parameter space of DDM
Table 2: Mean values and $1\,\sigma$ (68%) errors of the sum of neutrino masses $\Sigma m_{\nu}$ from different datasets in the $\Lambda\omega$DM, SDDM, DDM and DDME models. We quote $2\,\sigma$ (95%) upper limits for $\Sigma m_{\nu}$.
         Data          $\Sigma m_{\nu}$ [eV]
         $\Lambda\omega$DM          CD          $<0.168$
         CDS          $<0.162$
         SDDM          CD          $<0.120$
         CDS          $<0.0994$
         DDM          CD          $<0.164$
         CDS          $<0.145$
         DDME          CD          $<0.187$
         CDS          $<0.170$
@@ C. The effect of curvature on the dynamical dark sector
The cosmic curvature $\Omega_{K}$ plays an important role in depicting the background evolution and structure formation of the universe. Here we study its impact on the coexistence of DDM and DDE in the DDME scenario. Overall, we find that $\Omega_{K}$ hardly affect our conclusions. Interestingly, CMB gives $\omega_{dma}=0.164^{+0.093}_{-0.057}$, which is a $2.53\,\sigma$ evidence for the redshift evolution of DM (see Fig.2 and Tab.1). The addition of $\Omega_{K}$ increases the DDM evidence from $1.82\,\sigma$ to $2.53\,\sigma$. CD, CP and CDP provides a $1.93\,\sigma$, $2.40\,\sigma$ and $1.31\,\sigma$ evidence of DDM, respectively. Furthermore, we find that C, CD, CP and CDP exhibit a $\sim 2\,\sigma$, $2.37\,\sigma$, beyond $2\,\sigma$, and $2.40\,\sigma$ signal of DDE, respectively. It is worth noting that in the planes of $\omega_{dm0}-\omega_{0}$, $\omega_{dm0}-\omega_{a}$, $\omega_{dma}-\omega_{0}$ and $\omega_{dma}-\omega_{a}$, C, CD, CP and CDP still support a beyond $2\,\sigma$ evidence of the coexistence of DDM and DDE in the dark sector of the universe. Additionally, the $\sim 2\,\sigma$ preference of a closed universe by Planck CMB data could be alleviated when simultaneously considering the DDM and DDE.
@@
D. Relieving neutrino mass tension with DDM
Employing the degeneracy between DE and neutrino masses, the DESI collaboration demonstrate that CPL DDE can help relax the $2\,\sigma$ upper limit to a safe region DESI:2025ejh . Here we verify that evolving DM can also well resolve the neutrino mass tension between cosmological observations and terrestrial experiments (see Fig.3 and Tab.2). Using the combination of CMB and DESI data, we give, respectively, the $2\,\sigma$ upper bound $\Sigma m_{\nu}<0.168$, 0.120, 0.164 and 0.187 eV in the $\Lambda\omega$DM, SDDM, DDM and DDME models. If adding the Pantheon+ SN data, the bounds are slightly tightened to $\Sigma m_{\nu}<0.162$, 0.0994, 0.145 and 0.170 eV. This implies that there exists a strong degeneracy between the DM EoS and neutrino mass, similar to the case of DE.
E. Destiny+
In Wang:2025owe , based on the latest cosmological observations, we have proposed the new fate of the universe called the “Big Stall” in the CPL DDE model. In this work, when considering the coexistence of DDM and DDE, we propose a new density of the universe, i.e., the “Super Rip” in the DDME model. They both support that the universe is ultimately dominated by DM rather than DE. In Figs.4 and 5, we show the $1\,\sigma-5\,\sigma$ regions of background quantities $\Omega_{m}(z)$, $q(z)$, $E(z)$ and $\omega_{dm}(z)$ from different datasets in the DDM, SDDM and DDME scenarios. The details of their properties and the Super Rip have been specified in the main text.
F. Special example: constraints on $\omega_{dm0}$
@@
G. Reconstructing $\omega_{dm}(z)$ with Guassian process
@@
where the prime denotes the derivative with respect to the redshift $z$. The background quantities $D^{\prime}$ and $D^{\prime\prime}$ in Eq.(4) can be easily obtained using the GP reconstructions. In our analysis, for simplicity, we consider a flat universe, namely $\Omega_{k}=0$ and assume $H_{0}=73.04\pm 1.04$ km s-1 Mpc-1 Riess:2021jrx . We also take $\Omega_{m}=0.3153\pm 0.0073$ and $\Omega_{b}=0.04932\pm 0.00052$ Planck:2018vyg1 . For DESI BAO data, assuming $r_{d}=147.09\pm 0.26$ Mpc, we transform $D_{M}(z)/r_{d}$ and $D_{H}(z)/r_{d}$ to $D$ and $D^{\prime}$, respectively. For CC, we directly transform $H(z)$ to $D^{\prime}$, while transforming $D_{L}(z)$ from SN and $\mathcal{R}(z_{c})=1.7488\pm 0.0074$ Planck:2015bue ($z_{c}=1089.0$ is the redshift of recombination) from CMB to $D$. Same as Seikel:2012uu , we use the so-called Matérn ($\nu=9/2$) covariance function in GP reconstructions. Based on Eq.(4) and different kinds of datasets, we use a modified version of the package GaPP to carry our the GP reconstructions.
@@
H. Comparison of DESI DR1 and DR2
In Figs.8 and 9, we demonstrate the robustness of our conclusions and show the improvements of the constraining power of DESI DR2 BAO measurements relative to DESI DR1 data in the SDDM and DDME models. One can easily find that the constraints from DESI DR1 and DR2 data on evolution of DDM is robust and that the high-precision DESI DR2 BAO data significantly compresses the parameter space of DDE.